2. a) Find Ts(x), the third degree Taylor polynomial about x -0, for the function e2 b) Find a bound for the error in the interval [0, 1/2] 3. The following data is If all third order differences (not...

2.

a) Find Ts(x), the third degree Taylor polynomial about x -0, for the function e2

b) Find a bound for the error in the interval [0, 1/2]

3. The following data is If all third order differences (not divided differences) are 2, determine the coefficient of x in P(x). prepared for a polynomial P of unknown degree P(x) 2 1 4

I need help with both. Thank you.
Answer #1

Solution:2 (a) Let f(x) = er.

The general form of a Taylor expansion centered at a of a function xf (r) is

f) (a) n!

. where (n)(r is the nth derivative of f(x).

We have to find the third degree Taylor polynomial about x=0, for the function f(x) = er.

Therefore, substituting a=0 in (i), we have

7l fin)

f(3) (0) 31

31

Since f(x) = er,

Therefore

f'(x)=2e^{2x}, f''(x)=4e^{2x},f^{3}(x)=2^3e^{2x}=8e^{2x}

\Rightarrow f(0)=e^{0}=1,f'(0)=2e^{0}=2, f''(0)=2^2=4,f^{(3)}(0)=2^3=8

Therefore, (ii) becomes

e^{2x}=1+2x+\frac{x^{2}}{2!}4+\frac{8}{3!}x^{3}

\Rightarrow e^{2x}=1+2x+2x^2+\frac{4}{3}x^{3}

Therefore, the required third degree Taylor polynomial about x=0, for the function f(x) = er.

is given by P_{3}(x)=1+2x+2x^2+\frac{4}{3}x^{3}

b) Bound for the error in the interval [0, 1/2].

The error term, R_{n} is given by R_{n}=\frac{f^{(n)}(c)}{(n+1)!}x^{(n+1)} for some c E 0,1/2 .

Now \frac{d^{n}(e^{2x})}{dx^{n}}=2^{n}e^{2x} .

In the given case, n=3 and according to Lagrange's error term, we need n+1=3+1=4 derivative of

f(x) = er.

That is \frac{d^{4}(e^{2x})}{dx^{4}}=2^{4}e^{2x}=f^{(4)}(x) .

Therefore, the error is given by

R_{3}=\frac{f^{(4)}(c)}{4!}x^{4}=\frac{x^{4}}{4!}2^4e^{2c}for some c E 0,1/2 .

Now we want to find the largest possible value of this error on the interval [0,1/2]. Looking at the equation of R_{n} we see this will happen when the numerator is maximized and the denominator is minimized, i.e. x=1/2 and c=0. Thus an upper bound for |R_{n}| is:

|R_{n}|=\frac{(1/2)^{4}}{4!}2^4e^{0}=\frac{1}{24}.

Therefore upper bound for the error in the interval [0, 1/2] is 1/24.